OpenStudy (anonymous):

Find the exact value of arccos (sin (pi divided by six)). For full credit, explain your reasoning.

OpenStudy (anonymous):

I need help explaining. I have the answer, but I need to get this question correct!

zepdrix (zepdrix):

Lemme explain something really quick in regards to trig functions.

zepdrix (zepdrix):



zepdrix (zepdrix):

If I were to ask you for \(\Large\rm \cos\theta\), you would know that it's 3/5. But more specifically, when taking the trig function of an angle, we're `solving for a numerical value`, the ratio of sides. If I asked for \(\Large\rm \arccos\dfrac{3}{5}\), This gives us back our angle \(\Large\rm \theta\). So when you take the inverse trig function of something, you're `solving for an angle`.

zepdrix (zepdrix):

So with that in mind,\[\Large\rm \arccos\left(\sin \frac{\pi}{6}\right)\]What does sin(pi/6) give you?

OpenStudy (anonymous):

1/2

zepdrix (zepdrix):

\[\Large\rm \arccos\left(\frac{1}{2}\right)\]Mmmm ok good. Since we're taking the inverse trig function of this value, we're solving for an angle yes? \[\Large\rm \arccos\left(\frac{1}{2}\right)=\theta\]

zepdrix (zepdrix):

Do you remember how to switch between ... like cosine and arccosine? You just switch the arguments.\[\Large\rm \cos \theta=(stuff) \qquad\to\qquad \arccos(stuff)=\theta\]

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

So we can write our inverse function back in terms of cosine. \[\Large\rm \cos \theta=\frac{1}{2}\] Do you remember which special angle gives you 1/2?

OpenStudy (anonymous):

30

zepdrix (zepdrix):

I made a tiny boo boo I guess. Since we `started with` the inverse cosine function, we have to remember that our output is restricted to quadrants 1 and 2. Ummm I think it's 60 isn't it? It turns out that 300 also gives us 1/2 but that is in quadrant 4 so that one doesn't work for us.

OpenStudy (anonymous):

Are you talking about in the Unit Circle? Cause the sin (π/6) gives us 1/2 which is the y value of 30 degrees the Unit Circle

zepdrix (zepdrix):

Correct, but now we're dealing with cosine, not sine.

zepdrix (zepdrix):

The inner function was sin(30) which gave us 1/2. Now we're trying to figure out what angle when we take the cosine of it, gives us 1/2. It's not 30 again. :o

OpenStudy (anonymous):

It's 60. :p now that I check my Unit Circle

zepdrix (zepdrix):

Ah yes :) good!